### v2

# Spectral characterization and performance of SPT-SLIM on-chip filterbank spectrometers

Abstract
The South Pole Telescope Shirokoff Line Intensity Mapper (SPT-SLIM) experiment is a pathfinder for demonstrating the use of on-chip spectrometers for millimeter (mm) Line Intensity Mapping (LIM). We present spectral bandpass measurements of the SLIM spectrometer channels made on site using a Fourier Transform Spectrometer during SPT-SLIMs first deployment the 2024-2025 austral summer observing season. Through this we demonstrate a technique for measuring the narrow band passes of the SPT-SLIM filterbanks that improves beyond the intrinsic resolution of a Fourier Transform Spectrometer.
I Introduction
Line Intensity Mapping (LIM) uses imaging spectroscopy to trace emission lines as a function of redshift, producing three-dimensional spatial/spectral maps of the large scale structure (LSS) in the Universe without needing to spatially resolve individual sources. In comparison to galaxy surveys, LIM can access larger volumes of space and higher redshifts. LIM in the sub-millimeter/millimeter-wavelength band can provide constraints on cosmology, including expansion history [1] and the neutrino properties, effective number of neutrino species and the sum of the neutrino masses [2]. While it has significant potential, sub-millimeter/millimeter (sub-mm/mm) LIM is still maturing. Significant advancements in the sensitivity and scalability of large field-of-view mapping spectrometers are needed for sub-mm/mm LIM experiments to be competitive with and complementary to modern galaxy surveys.
SPT-SLIM is a pathfinder LIM experiment leveraging integrated on-chip filterbank spectrometers to place constrains on the CO power spectrum in the 120-180 GHz range from the SPT. SPT-SLIM targets CO(2-1) between $0.3<z<0.9$, CO(3-2) between $0.8<z<1.8$, and CO(4-3) between $1.6<z<2.8$ [3]. SPT-SLIM deployed deployed in the austral summer of 2024-2025 on the SPT in the camera position occupied by the Event Horizon Telescope (EHT) very-long-baseline interferometry (VLBI) receiver [4]. SPT is an ideal telescope for mm-wave LIM, as the high-altitude, low humidity, and stable weather at its location limit atmospheric noise [5]. The first deployment of SPT-SLIM has primarily focused on instrumentation and technology demonstration. Bright calibration sources, such as RCW-38 and the Moon, were prioritized to characterize the instrument. Future deployments will have updated detectors and will target a measurement of CO emission lines for LIM science.
To reconstruct the redshift dimension of the LIM signal, constrain atmospheric contamination, and disentangle interloper spectral features, reliable spectral response measurements of SPT-SLIM’s spectral channels are crucial. This work gives an overview of the SPT-SLIM detector design and details the spectral characterization of the SPT-SLIM spectrometers during the first SPT-SLIM deployment. We demonstrate measurements of the central frequencies and spectral bandpasses with suitable accuracy beyond the intrinsic resolution of the on-site Fourier Transform Spectrometer (FTS) via spatial-domain fitting of the FTS measurements [6, 7].
As a companion to this work, details on the SPT-SLIM Adiabatic Demagnetization Refrigerator (ADR) used in the deployment are given in Young et al. (2025) [8]. Additionally, details on the RF-ICE detector readout system can be found in Rouble et al. (2022) [9].
II SPT-SLIM: A Filterbank Spectrometer
In the SPT-SLIM design, light is coupled onto polarization sensitive orthomode transducers (OMTs) via gold-plated conical feedhorns [10, 11]. The OMTs are patterned on a thin film and mounted on an aluminum base with $\lambda/4$ backshort features. Each polarization mode of the OMT is coupled via a niobium microstrip transmission line to its own filterbank [12]. Each of the two filterbanks per antenna consist of 65 half-wave resonator filters capacitively coupled to the transmission line. Each filter provides a narrow spectral bandpass at a set frequency of light, with spectral resolution $R=\nu/\Delta\nu$, that then terminates on a Kinetic Inductance Detector (KID). SPT-SLIM uses a Lumped Element KID (LEKID) design with aluminum inductors and niobium interdigitated capacitors. All LEKIDs within a filterbank are coupled to a single microstrip readout line [13]. In addition to the narrow-band detectors that build up the spectrometer, each filterbank has three broadband detectors, and two dark detectors to assist in performance characterization. The dark detectors have no optical coupling to the filterbank transmission line. The SPT-SLIM focal plane consists of three wafers, called “submodules”, each with three readout lines, six filterbanks (one per polarization), and three antennae (see Fig. 1). Due to a combination of damage to the thin film OMTs and non-operational readout lines, five of the nine lines were operational for this first deployment. For additional details on the simulation, design, and fabrication of SPT-SLIM, see Barry et al. (2022) [12], Robson (2024) [13], Robson et al. (2022) [14], and Cecil et al (2023) [15].
SPT-SLIM adds to the burgeoning landscape of on-chip filterbank spectrometers. SPT-SLIM builds on the design and laboratory testing of SuperSpec [16] which has been successfully demonstrated through lab measurements [17] providing a multi-pixel on-sky demonstration of the technique in the 255–278 GHz range. SPT-SLIM also complements the single-pixel DEep Spectroscopic HIgh-redshift MApper (DESHIMA) experiment, operating on-sky in the 220–440 GHz range [18].
Figure 1: A schematic diagram of the full SPT-SLIM focal plane [13].
III Measurement
Figure 2: The spectral channels of a ploarisation-pair of deployment SPT-SLIM filterbanks measured with the FTS. Note that the width of each filter profile has been broadened by the instrumental line shape of the lower-resolution $\Delta\nu\sim 6$ GHz FTS.
Characterization of the bandpasses of the SPT-SLIM experiment requires both accurate determination of band centers and bandpass shape. An FTS is an interferometer capable of measuring the spectral response of the filterbank spectrometers with an excellent absolute frequency calibration, but with frequency resolution limited by the size of the FTS. A measurement is performed by moving a mirror on a motorized platform at a constant speed, varying the differential path length and introducing a phase difference between a stationary arm and the moving arm of the interferometer. The resulting interference signal (i.e., interferogram) measured by the detectors is the Fourier transform of the detector spectra. Here we note that spectroscopic wavenumbers $\sigma=\frac{1}{\lambda}=\frac{\nu}{c}$, where $c$ is the speed of light and $\nu$ is spectral frequency in units of inverse time, provides a natural frequency unit for FTS measurements. The focal plane was measured with the on-site FTS at the South Pole having a minimum resolution element of $d\nu\sim 6$ GHz, corresponding to $R=25$ at 150 GHz, limited by optical systematics [19]. Additionally, we note that the minimum resolvable spectral resolution element, $\Delta\nu$, is constant across the spectral band.
The SPT-SLIM spectrometers have a design resolution of $R\sim 100$, significantly greater than that of the FTS at the South Pole. As a result, the spectrometer channel bandpasses are unresolved and broadened by the instrumental line shape of the low resolution FTS. Measuring the bandpasses from the spectrum retrieved by the FTS would provide a biased measurement of the filterbank $R$ (see Fig. 2). We demonstrate that an accurate measurement of an under-resolved filter bandpass can be obtained by relying on the assumed spectral profile of the filters. A least-squares fit of the simulated detector signal measured in the spatial displacement domain of the FTS is used to extract the filter resolution and central frequency. For this, we have used the functional form of a Fourier-transformed Lorentzian (the corresponding profile of a half-wave filter),
$I(z)=e^{-\sigma_{0}\pi z/R}\cos{(2\pi\sigma_{0}z)},$ (1)
where $\sigma_{0}$ is the central spectral frequency of the band-pass in wavenumbers and $z$ is optical path difference.
Fig. 3 demonstrates this technique for an SPT-SLIM filter channel measured by the low-resolution FTS at the South Pole. The top panel shows that the interference signal is well fit to within the approximate uncertainty of the technique (see Fig. 4) in the spatial domain. In the Fourier domain, the spectral profile measured directly with the FTS is smeared out by the $\sim$6 GHz-wide line shape of the FTS while the Fourier transform of the functional form fitted in the spatial domain recovers a narrower, unbiased spectral profile (Fig 3, bottom panel).
[…]
Due to the interaction with other filters in the spectrometer, we expect the Lorentzian band-pass of a single filter to experience some asymmetry [13]. Despite the broadening from the FTS line profile, this can be seen in the measured spectral profile shown in Fig. 3 (lower panel). To account for this, we introduced a skew to the simulated data of the form
$S(\sigma)=\frac{2\tau}{1+2\pi i\tau(\sigma-\sigma_{0})}\left(1+\frac{\alpha}{\tau}(\sigma-\sigma_{0})\right),$ (2)
where $\tau=R/(\pi\sigma_{0})$ and $\alpha$ is the introduced skew parameter (see Fig. 5). For values ranging $|\alpha|=0$$10\tau$ with a SNR of 10, we have found that the asymmetry does not significantly affect the bias in values of $\sigma_{0}$ and $R$ extracted from spatial-domain fits.
Figure 3: A spatial-domain fit of the interference signal from one of the SLIM detectors (top). The spectrum of the corresponding spectrometer channel measured by the FTS and the spectral profile obtained from the spatial-domain fit are shown in the spectral (Fourier) domain (bottom). Note that the measured profile is smeared to a wider spectral bandpass due to the instrumental line shape of the low-resolution FTS.
Figure 4: The uncertainty in measurements of the central frequency (top) and spectral resolution/bandpass (bottom) of a filter extracted from spatial-domain fitting using simulated FTS measurements (see Eqs. 13). Simulated measurements were generated with a detector noise model at different signal-to-noise ratios (SNRs) and at different FTS resolving powers. In each case, the ratio of the resolving power of the South Pole FTS at 150 GHz to the $R=100$ simulated filter resolution is marked by a vertical dashed line.
Figure 5: The simulated spectral band pass of a filter with and without a modeled asymmetry (see Eq. 2).
IV Frequency Scheduling and Resolution
Figure 6: The fitted resolution, $R$, of the on-sky submodules extracted from measurements with the South Pole FTS (see Eq. 3). Error bars represent the biases determined from the simulated data shown in Fig. 4 at the corresponding SNR of the interferogram measured with the FTS.
Spatial-domain fitting was applied to the FTS measurements of both SPT-SLIM submodules that were operational for its first deployment at the SPT. All three lines on either submodule were measured, although one was not operational for on-sky measurements. These results are shown in Fig. 6. These measurements show a median spectral resolution of $33\pm 9$ for both deployment submodules, below the design target of $R\approx 100$. We have taken a conservative uncertainty estimate of $\pm 9$ in $R$, corresponding to the bias from the simulated results in Fig. 4 for the mean SNR of the filters in each submodule. Additionally, the centers of the bandpasses are shifted downward from the target SLIM band by $\sim$10 GHz.
Figure 7: A diagram of a single SPT-SLIM channel. Each filter (i.e., half-wave resonator) is capacitive coupled to the transmission line from the OMT and terminates in a capacitive coupled LEKID. Adapted from Robson (2024) [13].
Each spectral filter consists of a half-wave resonator connected to the transmission line and a detector via coupling capacitors which set the central frequency of the filter’s transmission (see Fig. 7). This was tuned primarily by the physical dimensions of the filter and coupling capacitors and informed by simulations employing projected fabrication errors and an anticipated dielectric loss tangent for SiN [14]. The shift to lower frequency in the SPT-SLIM bands suggests a systematic deviation from our simulated results that is under study for the next generation of SPT-SLIM spectrometers.
The resolution of a filter is governed by the joined quality factors of the filter’s coupling to the feedline $Q_{\mathrm{feed}}$, the detector coupling, $Q_{\mathrm{det}}$, and losses in the resonator $Q_{\mathrm{loss}}$,
$\frac{1}{R}=\frac{1}{Q_{\mathrm{filt}}}=\frac{1}{Q_{\mathrm{feed}}}+\frac{1}{Q_{\mathrm{det}}}+\frac{1}{Q_{loss}}.$ (3)
In SPT-SLIM, $Q_{\mathrm{feed}}$ and $Q_{\mathrm{det}}$ are largely controlled by the dimensions of the capacitors that couple the half-wave resonator to the feedline and to the detector, respectively. We expect $Q_{\mathrm{loss}}$ to be dominated by the dielectric loss tangent ($\tan\delta$) of the SiN employed in the microstrip architecture.
The coupling quality factors of SPT-SLIM were designed to provide an $R=100$ spectrometer with $\tan\delta=1\times 10^{-3}$. For these coupling quality factors, the resolution measurements in Fig. 6 would suggest a $\tan\delta\sim 2\times 10^{-2}$. This greatly exceeds what has been measured in characterization experiments of Argonne SiN and optical measurements of test structures on deployment SLIM submodule wafers [21, 13]. As such, the low resolution of the deployment SPT-SLIM submodules is likely a consequence of both poor dielectric performance and capacitor fabrication.
The dielectric loss tangent of the SPT-SLIM structures compromises its performance in nearly every aspect, reducing optical efficiency, resolution, and increasing optical loading from the atmospheric emission lines that bookend the SLIM band.
V Conclusion
In the 2024-2025 austral summer, SPT-SLIM deployed on the SPT in the camera position occupied by the EHT VLBI receiver. This deployment included 5 operational dual-polarization filterbank spectrometers. These spectrometers were characterized on site with a low-resolution FTS. We demonstrate a technique by which the spectral bandpass/resolution of the filters can be measured with an FTS having less than a quarter the resolution of the designed filter bandwidth. With this, the SPT-SLIM filterbanks deployed were measured to have an average spectral resolution $R=34$, lower than the target resolution of $R=100$. This is likely caused by differences between the simulated and fabricated dielectrics, in which lossy dielectric lowers the resolution from design. These measured bandpasses are applied to observations of the atmosphere and moon in Dibert et al. [22].
VI Acknowledgments
This work was supported by funding from the National Science Foundation, Fermilab National Laboratory under the Department of Energy, and the United Kingdom Research and Innovation.